SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report - March 12, 2003  Day 071 

Status Report From: China National Space Administration
Posted: Wednesday, March 12, 2003

DMSP 

DMSP operations were nominal over the past 24 hours. 

POES 

NOAA-11 rev 74579 / W at 1624Z on March 11:  Executed SCR #60, to conduct the first segment of SBUV end of life testing.  During this pass two commands were executed:  CXPIO 34D0 to select grating drive flex memory 1, and the SBCST strobe command to confirm flex memory switch.  All commanding was nominal. 

NOAA-11 rev 74581 / F at 1941Z on March 11:  Call-up pass to review SBUV telemetry, to ensure that no problems are evident as a result of the previous SBUV commanding, during rev 74579.  The instrument operational characteristics were unchanged since the last N-11 pass (i.e, the grating drive range did not change as commanded). 

NOAA-17 rev 3700 / F at 2111Z on March 11:  Two data sets from this support (G1A and G3A) lost frame synch lock during the course of the post-pass transfers from Fairbanks to CEMSCS.  The crew retried the transfers several different times before successfully getting all the data to SOCC and to CEMSCS.  The CDA is investigating suspected ground system issues, but the suspicion is that the problem is IMUX-related. 

NOAA-12 rev 61433 / W at 0953Z on March 12:  Executed command MSPOF to turn off the MSU scan motor as a result of the previous channel 1 and 2 drive belt failure.  This instrument was first declared operational on 9/17/91. 

POES operations were nominal over the past 24 hours. 

GOES 

At 1902Z on March 11, GOES-8 support was switched from COMM5 to COMM4 for maintenance purposes.  During the ensuing activities, the command comm controllers were inadvertently changed and the GOES-8 1915Z image was missed as a result. 

At 0402Z on March 12, a GOES-8 frame break was observed and attributed to loss of signal synchronization. 

Eclipse operations for all spacecraft continue.  Eclipse schedules for GOES-8 and GOES-10 (which can be used to track the day-to-day image deletions due to KOZ and eclipse) can be found at:  http://www.oso.noaa.gov/goes/koz-eclipse/koz-eclipse.htm 

GOES-12 remains in a test phase, in normal on-orbit mode, drifting at an eastward rate of ~0.33 degrees per day toward the GOES-East position where it will replace GOES-8 in April.  Routine SXI imaging is proceeding along with SXI test exposures.  SXI testing took place yesterday between 1458Z and 1647Z and between 1913Z and 1925Z. 

GOES-9 is currently at approximately 172.8 W drifting at a westward rate of ~0.75 degrees per day toward 155 E longitude where it will support the GMS-5 replacement mission in April.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter